### v1

@@ 3.1 Simulation Description
In this section, we analyze these systems numerically. We do so using the publicly available code ${\mathcal{C}}$${\mathcal{L}}$attice Figueroa:2020rrl ; Figueroa:2021yhd to simulate a scalar field with an exponential potential in a flat FLRW universe (to be clear, ${\mathcal{C}}$${\mathcal{L}}$attice does a full numerical simulation of the scalar field but does not simulate the metric; it does not do numerical GR). A full numerical GR simulation of a kinating scalar field will appear shortly in KingsGroup .
We use a $32^{3}$ sized lattice, which is sufficient for us to keep our errors under $0.1\%$ at all times. In these simulations, the Hubble constant is given in terms of the average energy density in the simulation box. For the simulation, ${\mathcal{C}}$${\mathcal{L}}$attice transforms the physical spacetime variables $\{\phi,dt,dx\}$ into dimensionless program variables $\{\tilde{\phi},d\tilde{\eta},d\tilde{x}\}$ through
@@
$k=\frac{2\pi N_{\lambda}}{L}=k_{IR}N_{\lambda}.$ (3.42)
Here $L=2\pi/k_{IR}$ is the physical size of the box being simulated, which is converted into the corresponding infrared cutoff $k_{IR}$. Meanwhile, $N_{\lambda}$ is the number of standing waves present in the box. The smallest relevant physical timescale is the oscillation period of the scalar field perturbations, which at time $t$ is given by [^5] This is found by expanding Eq. (3.41) around $t\rightarrow t+\Delta t$ for $\Delta t\ll t$, noting that $a^{-1}t\propto t^{1-c}$.
@@
To resolve each oscillation with the same number of program time steps we need $d\eta$ to scale like $\Delta t$, which requires setting $\alpha=1$. Since e-folds scale with time as $dN\propto a^{1-1/c}d\eta$, this implies that each successive e-fold takes longer to simulate than the previous one in cosmologies with $c<1$. To ensure that perturbations behave like radiation in non-kination epochs, we want them to be sub-horizon. We therefore set the initial ratio of the comoving wavelength $\lambda$ to the comoving Hubble length $r_{H}=(aH)^{-1}$ to be comparable to unity $(\lambda r_{H}^{-1})_{i}\lesssim 1$. In cosmologies with $c<1$, with time they become more sub-horizon since the Hubble horizon grows in comoving space
$\lambda r_{H}^{-1}=(\lambda r_{H}^{-1})_{i}a^{1-1/c}.$ (3.44)
@@
$E_{g,i}=\frac{1}{2}\sum_{j=1}^{3}\langle(\partial_{j}\phi)^{2}\rangle=\frac{3}{4}A_{i}^{2}N_{\lambda}^{2}k_{IR}^{2},$ (3.45)
@@
$x^{2}_{i}=\frac{\bar{\pi}_{i}^{2}}{6H_{i}^{2}M_{P}^{2}},\ \ \ \ \ \ \ y^{2}_{i}=\frac{V_{i}e^{-\alpha\bar{\phi}(t_{i})/M_{P}}}{3H_{i}^{2}M_{P}^{2}}.$ (3.47)
Desired initial conditions can then be acquired by picking appropriate $A_{i}$, $\bar{\pi}_{i}$, and $\bar{\phi}(t_{i})$ and an initial Hubble constant $H_{i}$.
@@
$z^{2}=\bigg{(}\frac{E_{k}}{3H^{2}M_{P}^{2}}-x^{2}\bigg{)}+\frac{E_{g}}{3H^{2}M_{P}^{2}}+\frac{\alpha^{2}\bar{V}\langle\delta\phi^{2}\rangle}{6H^{2}M_{P}^{4}}.$ (3.50)
@@
$y^{2}=\frac{E_{p}}{3H^{2}M_{P}^{2}}-\frac{\alpha^{2}\bar{V}\langle\delta\phi^{2}\rangle}{6H^{2}M_{P}^{4}}.$ (3.51)
@@ 3.2 Simulation Results
Figure $(\lambda r_{H}^{-1})_{i}$ $N_{\lambda}$ $x^{2}_{i}$ $y^{2}_{i}$ $z^{2}_{i}$
1 1.4 2 0.495 0.498 0.007
2 0.17 2 0.23 0.16 0.61
@@
$\tilde{k}^{2}=\frac{15}{4}-\frac{16}{\alpha^{2}}.$ (3.56)
@@ 4.1 Duration of Kination
where $\bar{\rho}=\bar{\pi}^{2}/2$ is the kination energy density. Meanwhile, $\delta=2\delta\pi/\bar{\pi}$ is the leading contribution, leading in the perturbation, which vanishes upon a spatial averaging, leaving behind $\Omega_{\delta}(k)=\tfrac{1}{2}\langle\delta^{2}\rangle$. This can be expressed in terms of the dimensionless power spectrum sourced by modes entering the horizon, from the time when kination begins ($k_{\text{kin}}$) until a later time ($k=aH$) [^7] The power spectrum is that for the linear perturbation $\mathcal{P}_{\delta}(k)=\int d^{3}\boldsymbol{x}\langle\delta(0)\delta(\boldsymbol{x})\rangle e^{i\boldsymbol{k}\cdot\boldsymbol{x}}$.
$\Omega_{\delta}(k)=\frac{1}{2}\int_{k}^{k_{kin}}d\ln\tilde{k}\ \mathcal{P}_{\mathcal{\delta}}(\tilde{k}).$ (4.59)
@@
$V(\phi)=V_{0}e^{-\alpha\phi/M_{P}}=V_{\text{kin}}e^{-\sqrt{6}\alpha\Delta N},$ (4.64)
@@
$\Delta N=\frac{1}{\alpha\sqrt{6}}\ln\bigg{(}\frac{V_{\text{kin}}}{V_{\text{IR}}}\bigg{)}.$ (4.65)
Kination occurs on exponential potentials as long as $\alpha>\sqrt{6}$ Copeland:1997et in which case for $V_{\text{IR}}=10^{-20}V_{\text{kin}}$, we find that $\Delta N\leq 7.7$. In the case of the Large Volume Scenario Balasubramanian:2005zx where $\alpha=\sqrt{27/2}$, the duration before reaching the minimum is $\Delta N=5.1$. Therefore, if the scalar field perturbations are sourced entirely from slow-roll inflation, then the resulting perturbation amplitude is insufficient for the post-inflationary scenario on exponential potentials to reach the self-tracker before the scalar field reaches the minimum.
@@ 4.2 Conclusion
In this paper we have shown that such a scenario can indeed be realized. Through both analytical computation and numerical simulation using ${\mathcal{C}}$${\mathcal{L}}$attice , we have found that the self-perturbations of the rolling scalar field can themselves serve as a ‘background’ radiation fluid. The numerical computations, which involve a full 3d simulation of the field, show that the results can be interpreted as a ‘background’ averaged value of the field evolving in the presence of a ‘radiation fluid’ sourced by the self-perturbations of the scalar field. The resulting fixed point of this solution has the energy proportions associated to conventional radiation trackers.
@@ Acknowledgements.
We thank Fien Apers, Ethan Carragher, Cheng Cheng, Panagiotis Giannadakis, Lucien Huertier, Eugene Lim and Noelia Sanchez-Gonzalez for discussions related to this topic. JC acknowledges support from the STFC consolidated grants ST/T000864/1 and ST/X000761/1 and is also a member of the COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology). MM is supported by the St John’s College Graduate Scholarship in partnership with STFC. EJC acknowledges support from the STFC Consolidated Grant ST/X000672/1. For the purpose of Open Access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.

### v2

@@ 3.1 Simulation Description
[…]
@@
$k=\frac{2\pi N_{\tilde{\lambda}}}{L}=k_{IR}N_{\tilde{\lambda}}$. (3.42)
Here $L=2\pi/k_{IR}$ is the physical size of the box being simulated, which is converted into the corresponding infrared cutoff $k_{IR}$. Meanwhile, $N_{\tilde{\lambda}}$ is the number of standing waves present in the box. The smallest relevant physical timescale is the oscillation period of the scalar field perturbations, which at time $t$ is given by [^6] This is found by expanding Eq. (3.41) around $t\rightarrow t+\Delta t$ for $\Delta t\ll t$, noting that $a^{-1}t\propto t^{1-c}$.
@@
To resolve each oscillation with the same number of program time steps we need $d\eta$ to scale like $\Delta t$, which requires setting $\alpha=1$. Since e-folds scale with time as $dN\propto a^{1-1/c}d\eta$, this implies that each successive e-fold takes longer to simulate than the previous one in cosmologies with $c<1$. To ensure that perturbations behave like radiation in non-kination epochs, we want them to be sub-horizon. We therefore set the initial ratio of the comoving wavelength $\tilde{\lambda}$ to the comoving Hubble length $r_{H}=(aH)^{-1}$ to be comparable to unity $(\tilde{\lambda}r_{H}^{-1})_{i}\lesssim 1$. In cosmologies with $c<1$, with time they become more sub-horizon since the Hubble horizon grows in comoving space
$\tilde{\lambda}r_{H}^{-1}=(\tilde{\lambda}r_{H}^{-1})_{i}a^{1-1/c}.$ (3.44)
@@
$E_{g,i}=\frac{1}{2}\sum_{j=1}^{3}\langle(\partial_{j}\phi)^{2}\rangle=\frac{3}{4}A_{i}^{2}N_{\tilde{\lambda}}^{2}k_{IR}^{2},$ (3.45)
@@
$x^{2}_{i}=\frac{\bar{\pi}_{i}^{2}}{6H_{i}^{2}M_{P}^{2}},\ \ \ \ \ \ \ y^{2}_{i}=\frac{V_{i}e^{-\lambda\bar{\phi}(t_{i})/M_{P}}}{3H_{i}^{2}M_{P}^{2}}.$ (3.47)
Desired initial conditions can then be acquired by picking appropriate $A_{i}$, $\bar{\pi}_{i}$, and $\bar{\phi}(t_{i})$ for some initial Hubble constant $H_{i}$.
@@
$z^{2}=\bigg{(}\frac{E_{k}}{3H^{2}M_{P}^{2}}-x^{2}\bigg{)}+\frac{E_{g}}{3H^{2}M_{P}^{2}}+\frac{\lambda^{2}\bar{V}\langle\delta\phi^{2}\rangle}{6H^{2}M_{P}^{4}}.$ (3.50)
@@
$y^{2}=\frac{E_{p}}{3H^{2}M_{P}^{2}}-\frac{\lambda^{2}\bar{V}\langle\delta\phi^{2}\rangle}{6H^{2}M_{P}^{4}}.$ (3.51)
@@ 3.2 Simulation Results
Figure $\lambda$ $N_{\tilde{\lambda}}$ $H_{i}$ (GeV) $(\tilde{\lambda}r_{H}^{-1})_{i}$ $x^{2}_{i}$ $y^{2}_{i}$ $z^{2}_{i}$
11 33 22 $4.5\times 10^{8}$ $1.4$ $0.495$ $0.498$ $0.007$
22 33 22 $5.5\times 10^{7}$ $0.17$ $0.23$ $0.16$ $0.61$
@@
$\tilde{k}^{2}=\frac{15}{4}-\frac{16}{\lambda^{2}}.$ (3.56)
@@ 4.1 Duration of Kination
where $\bar{\rho}=\bar{\pi}^{2}/2$ is the kination energy density. Meanwhile, $\delta=2\delta\pi/\bar{\pi}$ is the leading contribution, leading in the perturbation, which vanishes upon a spatial averaging, leaving behind $\Omega_{\delta}(k)=\tfrac{1}{2}\langle\delta^{2}\rangle$. This can be expressed in terms of the dimensionless power spectrum sourced by modes entering the horizon, from the time when kination begins ($k_{\text{kin}}$) until a later time (k=aHk=aH) [^8] The power spectrum is that for the linear perturbation $\mathcal{P}_{\delta}(k)=\int d^{3}\boldsymbol{x}\langle\delta(0)\delta(\boldsymbol{x})\rangle e^{i\boldsymbol{k}\cdot\boldsymbol{x}}$.
[…]
@@
$V(\phi)=V_{0}e^{-\lambda\phi/M_{P}}=V_{\text{kin}}e^{-\sqrt{6}\lambda\Delta N},$ (4.64)
@@
$\Delta N=\frac{1}{\lambda\sqrt{6}}\ln\bigg{(}\frac{V_{\text{kin}}}{V_{\text{IR}}}\bigg{)}.$ (4.65)
Kination occurs on exponential potentials as long as $\lambda>\sqrt{6}$ Copeland:1997et in which case for $V_{\text{IR}}=10^{-20}V_{\text{kin}}$, we find that $\Delta N\leq 7.7$. In the case of the Large Volume Scenario Balasubramanian:2005zx where $\lambda=\sqrt{27/2}$, the duration before reaching the minimum is $\Delta N=5.1$. Therefore, if the scalar field perturbations are sourced entirely from slow-roll inflation, then the resulting perturbation amplitude is insufficient for the post-inflationary scenario on exponential potentials to reach the self-tracker before the scalar field reaches the minimum.
@@ 4.2 Conclusion
[…]
@@ Acknowledgements.
We thank Fien Apers, Ethan Carragher, Cheng Cheng, Panagiotis Giannadakis, Lucien Heurtier, Eugene Lim and Noelia Sanchez-Gonzalez for discussions related to this topic. JC acknowledges support from the STFC consolidated grants ST/T000864/1 and ST/X000761/1 and is also a member of the COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology). MM is supported by the St John’s College Graduate Scholarship in partnership with STFC. EJC acknowledges support from the STFC Consolidated Grant ST/X000672/1. For the purpose of Open Access, the authors have applied a CC BY public copyright licence to any Author Accepted Manuscript version arising from this submission.